# First-Mover Bias in Gradient Boosting Explanations: Mechanism, Detection, and Resolution

### 22 Mar 2026 v1

@@ Abstract
We provide evidence that model independence is sufficient to largely neutralize first-mover bias in the linear, high-collinearity regime we study, and that it remains the most effective mitigation under nonlinear data-generating processes. Both our proposed method, DASH (Diversified Aggregation of SHAP), and simple seed-averaging (Stochastic Retrain) restore stability by breaking the sequential dependency chain, supporting the view that the operative mechanism is independence between explained models, not any particular aggregation strategy. At $\rho=0.9$, both methods achieve stability =0.977=0.977, while the standard single-best workflow degrades to $0.958$ and the Large Single Model to $0.938$. On the Breast Cancer dataset, DASH improves stability from $0.32$ to $0.93$ (+0.61+0.61) over the training-budget-matched Single Best (M=200M{=}200).
DASH additionally provides two novel diagnostic tools—the Feature Stability Index (FSI) and Importance-Stability (IS) Plot—that detect first-mover bias without ground truth and can be applied independently of the full DASH pipeline, enabling practitioners to audit explanation reliability before acting on feature rankings. Software and reproducible benchmarks are available at https://github.com/DrakeCaraker/dash-shap.
@@ 1 Introduction
The intuitive response to unstable explanations is to train a more powerful model. Our results suggest that this can be counterproductive. A single large XGBoost model with thousands of trees and the same total tree count as DASH ($M\times{\sim}75\approx 15{,}000$ trees, where ${\sim}75$ is the average number of trees per population model after early stopping) produces the poorest attribution reproducibility among all methods tested. We hypothesize, and our experiments support, that the primary mechanism is sequential residual dependency: in gradient boosting, each tree fits the residuals of all previous trees. If tree 1 selects feature $A$ from a correlated pair $(A,B)$, it partially removes $A$’s signal from the residuals. Subsequent trees find both $A$ and $B$ less useful, but $A$ slightly more so because the residual structure favors the feature that was partially captured. Over thousands of iterations, this creates a “first-mover advantage” that concentrates importance on whichever feature happened to be selected first—an artifact of optimization path dependence, not a property of the data. We use the term first-mover bias to denote this specific path-dependent concentration effect in sequential boosting; our claim is not that this is the only source of SHAP instability, but that it is an important and previously under-emphasized contributor in gradient-boosted trees under correlation.
We note that with low colsample_bytree (0.1–0.5 in our population), correlated features $A$ and $B$ are not always present in the same column sample, so the first-mover advantage is probabilistic rather than deterministic: over many trees, whichever feature accumulates slightly more early selections gains a cumulative residual advantage that is attenuated—but not eliminated—by column subsampling.
@@
We make four contributions:
@@
  • Infrastructure. The fit_from_attributions() interface decouples the DASH aggregation stage from XGBoost, making the pipeline applicable to any attribution method (LIME, Integrated Gradients, neural network SHAP) that produces feature-level attribution vectors. This positions DASH as a general-purpose stability layer for model explanation rather than an XGBoost-specific tool.
@@ 2 Related Work
The instability of SHAP values under model perturbation has been noted by several authors. Fisher et al. (2019) formalize the Rashomon set—the collection of models with near-optimal performance—and show that variable importance can vary widely across this set. Dong and Rudin (2020) visualize the “cloud” of variable importance across Rashomon-set models, demonstrating that importance rankings are inherently ambiguous when many near-optimal models exist. Semenova et al. (2022) further characterize the Rashomon set’s structure and its implications for model selection. Marx et al. (2023) demonstrate that SHAP-based feature attributions are sensitive to reference distribution choices. Covert et al. (2021) provide a unified framework connecting removal-based explanations (including SHAP) and discuss stability considerations across methods. The problem is especially acute for correlated features, where SHAP must distribute credit among features that carry overlapping information (Kumar et al., 2020; Chen et al., 2020).
@@
Paillard et al. (2025) address a distinct consistency problem: given a single already-fitted ensemble model, they argue for computing one SHAP attribution over the full ensemble rather than averaging per-tree attributions, for internal consistency within that model. Their setting does not address cross-fit reproducibility—whether retraining the model on the same data yields the same attributions. Our results show that this cross-fit instability (our focus) is not resolved by the single-ensemble approach: the single-ensemble baseline (“Ensemble SHAP”) does not improve stability over Single Best, and the Large Single Model, which maximizes sequential tree dependency, produces the worst stability of any method tested.
@@ 3.3 Sequential residual dependency
We hypothesize that for a gradient-boosted model $f=\sum_{t=1}^{T}h_{t}$ with $T$ trees, if features $j,k$ belong to a correlated group with pairwise correlation $\rho\to 1$ and tree $h_{1}$ splits on feature $j$, then $\mathbb{E}[|\phi_{j}(f)|]>\mathbb{E}[|\phi_{k}(f)|]$ under TreeSHAP, with the gap increasing in $T$. The expectation is over the randomness in data sampling and split selection. We test the $T$-dependence indirectly in Section 6.1 via the Large Single Model comparison: the LSM uses ${\sim}15{,}000$ sequential trees and produces the poorest reproducibility of any method, consistent with the prediction that more sequential iterations amplify first-mover concentration. A direct isolation experiment varying $T$ while holding other factors constant is available in the code repository and shows monotonically increasing concentration with tree count.
@@ 4.2 Stage 2: Performance Filtering
We retain models whose validation score is within $\varepsilon$ of the best. Two modes are supported:
$\mathcal{F}_{\text{abs}}=\{i:|s_{i}-s^{*}|\leq\varepsilon\},\qquad\mathcal{F}_{\text{rel}}=\{i:|s_{i}-s^{*}|\leq\varepsilon\cdot|s^{*}|\},\quad s^{*}=\max_{i}s_{i},$ (2)
where $s_{i}$ is the negative RMSE (for regression) or AUC (for classification) of model $i$. This ensures that all explanations come from models that have learned meaningful signal. We use $\varepsilon=0.08$ (absolute mode, $\mathcal{F}_{\text{abs}}$) for synthetic data and $\varepsilon=0.05$ (relative mode, $\mathcal{F}_{\text{rel}}$, retaining models within 5% of the best score) for real-world datasets.
@@ 4.3 Stage 3: Diversity Selection
This ensures each selected model is maximally different from all previously selected models in its feature utilization pattern, without requiring knowledge of the feature correlation structure.
@@ 5.1 Synthetic data
We sweep $\rho\in\{0.0,0.5,0.7,0.9,0.95\}$ with $N_{\text{reps}}=20$ repetitions at each level, regenerating data (same coefficients, new random draws) for each repetition.
@@ 5.2 Real-world datasets
For synthetic data, each of the 20 repetitions regenerates the dataset (same coefficients, new random draws) and retrains all models, capturing both data-sampling and model-selection variance. For real-world datasets, each repetition retrains all models with different random seeds on the same fixed data split, isolating model-selection variance from data-sampling variance. This distinction means that real-world stability estimates reflect only the instability due to arbitrary model choices, not data variability.
@@ 5.3 Methods compared
Independent Stochastic Retrain $K$ models, different seeds, single fixed hyperparameter config (best of 100 random draws from the same grid as DASH)
@@ 5.4 Evaluation metrics
Mean pairwise Spearman correlation across $N_{\text{reps}}=20$ repeated runs of each method:
@@ 6.1 The mechanism: evidence for first-mover bias
Table 3 presents results for the four principal methods across $\rho\in\{0.0,0.5,0.7,0.9,0.95\}$ with 20 repetitions per level. The remaining five baselines (Ensemble SHAP, Single Best M=200M{=}200, LSM Tuned, Random Selection, Naive Top-$N$) are evaluated at $\rho=0.9$ only (Table 4), as their primary purpose is to contextualize the mechanism at high correlation. The key finding is that LSM achieves the poorest stability of any method at every correlation level, despite matching DASH’s total tree count and feature restriction. This is consistent with the first-mover bias hypothesis: sequential residual dependency concentrates importance on arbitrary features, and the effect worsens with correlation severity (LSM stability degrades from $0.953$ at $\rho=0$ to $0.925$ at $\rho=0.95$). Meanwhile, DASH’s stability is effectively flat (0.9720.9720.9770.977), demonstrating immunity to the mechanism it is designed to break.
Table 3: The mechanism experiment: DASH vs. Single Best vs. Large Single Model across correlation levels (20 repetitions per $\rho$). Both DASH and LSM use the same low colsample_bytree; the critical contrast for our hypothesis is that DASH’s models are trained independently while LSM trains trees sequentially. Bold indicates best per metric per $\rho$ level. The four principal methods are shown; all nine are compared at $\rho=0.9$ in Table 4.
@@
0.0 Single Best $973\pm.001 .$985 .168 .609
Large Single Model $953\pm.003 .$974 .170 .771
Stoch. Retrain $\mathbf{.975}\pm.001$ .987 .175 .581
DASH (MaxMin) $972\pm.002 .$985 .164 .596
0.5 Single Best $975\pm.002 .$987 .178 .617
Large Single Model $965\pm.002 .$981 .194 .767
Stoch. Retrain $\mathbf{.980}\pm.001$ .989 .173 .585
DASH (MaxMin) $977\pm.001 .$988 .163 .599
0.7 Single Best $969\pm.002 .$984 .193 .618
Large Single Model $963\pm.002 .$981 .209 .757
Stoch. Retrain $\mathbf{.980}\pm.001$ .990 .170 .582
DASH (MaxMin) $977\pm.001 .$989 .162 .597
0.9 Single Best $958\pm.003 .$978 .224 .614
Large Single Model $938\pm.003 .$967 .262 .738
Stoch. Retrain $\mathbf{.977}\pm.002$ .988 .182 .577
DASH (MaxMin) $\mathbf{.977}\pm.001$ .988 .176 .594
0.95 Single Best $951\pm.004 .$975 .246 .608
Large Single Model $925\pm.003 .$961 .284 .733
Stoch. Retrain $\mathbf{.979}\pm.002$ .989 .170 .576
DASH (MaxMin) $977\pm.001 .$988 .172 .590
@@
    The effect scales with correlation. LSM stability degrades from $0.953$ to $0.925$ as $\rho$ increases from 0 to $0.95$. DASH stability is flat (0.9720.9720.9770.977). First-mover bias is specifically a collinearity problem.
@@
    Equity degrades in the same pattern. LSM’s within-group CV worsens from $0.170$ to $0.284$, consistent with sequential dependency concentrating importance within correlated groups. DASH distributes it proportionally (0.1640.1640.1760.176).
@@ 6.2 The principle: independence largely resolves it
Table 4 compares all methods at $\rho=0.9$. The critical observation is that DASH and Stochastic Retrain achieve identical stability ($0.977$ each). These methods differ substantially in design— DASH uses forced feature restriction, performance filtering, and diversity-aware selection, while Stochastic Retrain trains $K$ models with different random seeds and a single fixed hyperparameter configuration (the best of 100 random draws from the same grid)—but they share one property: their explained models are trained independently.
Table 4: All methods at $\rho=0.9$ (20 repetitions), grouped by whether they achieve model independence. Bold indicates best per metric. Matched training budget (M=200M{=}200 models trained). Timing entries marked — share infrastructure with other methods and were not independently measured. 95% BCa bootstrap CIs on stability: Single Best $[0.952,0.963]$, DASH $[0.975,0.978]$ (non-overlapping). Stochastic Retrain: $0.977\pm 0.0015$ (SE); BCa interval overlaps with DASH ($0.977\pm 0.0012$ SE), consistent with equivalence.
@@
Dependent Single Best $958\pm.003 .$978 .224 043.5
Single Best $964\pm.002 .$981 .212 248.8
Large Single Model $938\pm.003 .$967 .262 006.4
LSM (Tuned) $948\pm.003 .$972 .272 088.9
Ensemble SHAP $956\pm.003 .$977 .237
Independent Stochastic Retrain $977\pm.002 .$988 .182 233.5
Random Selection $976\pm.001 .$988 .187 287.2
Naive Top-$N$ $976\pm.001 .$988 .187
DASH (MaxMin) $\mathbf{.977}\pm.001$ .988 .176 140.3
@@
The gap between tiers ($\sim$0.01–0.04) dwarfs the gap within the independent tier ($\sim$0.001). This supports the view that the operative mechanism is independence between explained models, not any particular aggregation strategy or hyperparameter diversification.
@@
Within the independent tier, Random Selection ($0.976$) nearly matches DASH ($0.977$), suggesting that MaxMin diversity selection contributes minimally to stability beyond what random sampling from the filtered population already provides. The primary value of diversity selection is therefore not stability but equity: DASH’s within-group CV ($0.176$) is lower than Random Selection’s ($0.187$), indicating that deliberate diversity in feature utilization patterns produces more equitable credit distribution (though this comparison has not been formally significance-tested). Practitioners who need only stability can use random selection from a performance-filtered population; those who also require equitable attributions benefit from the full DASH pipeline.
@@
To directly quantify the role of model-selection randomness, we decompose explanation instability into two sources by alternately fixing the data seed or the model seed. When data is fixed and only model seeds vary (isolating model-selection variance), Single Best stability is $0.978$ while DASH achieves 0.9950.995—indicating that DASH nearly eliminates model-selection noise. Model-selection variance accounts for approximately $54\%$ of Single Best’s total instability but only $24\%$ of DASH’s, consistent with independence-based averaging cancelling the arbitrary feature choices that dominate single-model explanations. [^2] These percentages use $1-\text{stability}$ as a proxy for instability. Since stability is a mean pairwise Spearman correlation rather than a variance, the decomposition is approximate and should be interpreted as directional evidence rather than exact variance fractions.
@@
Table 5 reports Wilcoxon signed-rank tests with Holm–Bonferroni step-down correction on accuracy and equity metrics across all $\rho$ levels. All comparisons between DASH and LSM are significant at every $\rho$ level with large effect sizes (Cohen’s $d>1.4$). DASH vs. Single Best becomes significant at $\rho\geq 0.7$ for both accuracy ($d=+0.98$) and equity (d=1.07d=-1.07). The DASH vs. Stochastic Retrain comparison is not significant on either accuracy ($d=+0.05$) or equity (d=0.21d=-0.21), consistent with the independence principle.
The small effect sizes ($d=+0.05$ for accuracy, d=0.21d=-0.21 for equity) and large Wilcoxon $p$-values ($p=0.926$ for accuracy, $p=0.622$ for equity) are consistent with practical equivalence: differences of this magnitude—well under $0.005$ in absolute accuracy—do not change top-5 feature rankings in practice. Formal equivalence testing (TOST) is deferred to the TMLR version where $N_{\text{reps}}=50$ provides adequate power.
Note that stability, computed as a single aggregate across all repetition pairs, cannot be subjected to per-repetition Wilcoxon tests; we report BCa bootstrap confidence intervals (Table 4 caption) as an alternative measure of precision for this metric.
Table 5: Wilcoxon signed-rank tests with Holm–Bonferroni step-down correction (20 paired repetitions). Bold $p$-values are significant at adjusted $\alpha=0.05$. Selected comparisons shown from 26 total tests; full results available in the code repository.

### 06 Apr 2026 v2

@@ Abstract
We provide evidence that model independence is sufficient to largely neutralize first-mover bias in the linear, high-collinearity regime we study, and that it remains the most effective mitigation under nonlinear data-generating processes. Both our proposed method, DASH (Diversified Aggregation of SHAP), and simple seed-averaging (Stochastic Retrain) restore stability by breaking the sequential dependency chain, supporting the view that the operative mechanism is independence between explained models, not any particular aggregation strategy. At $\rho=0.9$, both methods achieve stability $\approx 0.977$$0.978$, while the standard single-best workflow degrades to $0.958$ and the Large Single Model to $0.938$. On the Breast Cancer dataset, DASH improves stability from $0.376$ to $0.925$ (+0.549+0.549) over Single Best, and from $0.339$ to $0.925$ (+0.586+0.586) over the training-budget-matched Single Best (M=200M{=}200), outperforming even Stochastic Retrain by +0.063+0.063. Under nonlinear data-generating processes, the stability advantage emerges at $\rho\geq 0.7$; at lower correlation, simpler methods suffice.
DASH additionally provides two diagnostic tools—the Feature Stability Index (FSI) and Importance-Stability (IS) Plot—that detect first-mover bias without ground truth and can be applied independently of the full DASH pipeline, enabling practitioners to audit explanation reliability before acting on feature rankings. Software and reproducible benchmarks are available at https://github.com/DrakeCaraker/dash-shap.
@@ 1 Introduction
The intuitive response to unstable explanations is to train a more powerful model. Our results suggest that this can be counterproductive. A single large XGBoost model with thousands of trees and the same total tree count as DASH ($M\times{\sim}75\approx 15{,}000$ trees, where ${\sim}75$ is the average number of trees per population model after early stopping) produces the poorest attribution reproducibility among all methods tested. We hypothesize, and our experiments support, that the primary mechanism is sequential residual dependency: in gradient boosting, each tree fits the residuals of all previous trees. If tree 1 selects feature $A$ from a correlated pair $(A,B)$, it partially removes $A$’s signal from the residuals. Subsequent trees find both $A$ and $B$ less useful, but $A$ slightly more so because the residual structure favors the feature that was partially captured. Over thousands of iterations, this creates a “first-mover advantage” that concentrates importance on whichever feature happened to be selected first—an artifact of optimization path dependence, not a property of the data. We use the term first-mover bias to denote this specific path-dependent concentration effect in sequential boosting; our claim is not that this is the only source of SHAP instability, but that it is an important and previously under-emphasized contributor in gradient-boosted trees under correlation. We note that with low colsample_bytree (0.1–0.5 in our population), correlated features $A$ and $B$ are not always present in the same column sample, so the first-mover advantage is probabilistic rather than deterministic: over many trees, whichever feature accumulates slightly more early selections gains a cumulative residual advantage that is attenuated—but not eliminated—by column subsampling.
@@
DASH operates at two layers. The core operation—averaging attributions across independently trained models—is proved to be the minimum-variance unbiased estimator (Cramér–Rao bound) and is Pareto-optimal among all stable attribution methods (Caraker et al., 2026b). This core operation requires only model independence; even simple seed averaging (Stochastic Retrain) implements it. The pipeline (Stages 1–5) adds diversity enforcement, performance filtering, and stability diagnostics that improve finite-sample equity and provide actionable audit output beyond what plain averaging offers.
We make five contributions:
@@
  • Infrastructure. The fit_from_attributions() interface decouples the DASH aggregation stage from XGBoost, making the pipeline applicable to any attribution method (LIME, Integrated Gradients, neural network SHAP) that produces feature-level attribution vectors. This positions DASH as a general-purpose stability layer for model explanation rather than an XGBoost-specific tool (validated on LIME attributions in Appendix G).
@@ 2 Related Work
The instability of SHAP values under model perturbation has been noted by several authors. Fisher et al. (2019) formalize the Rashomon set—the collection of models with near-optimal performance—and show that variable importance can vary widely across this set. Dong and Rudin (2020) visualize the “cloud” of variable importance across Rashomon-set models, demonstrating that importance rankings are inherently ambiguous when many near-optimal models exist. Semenova et al. (2022) further characterize the Rashomon set’s structure and its implications for model selection. Bilodeau et al. (2024) prove impossibility results for complete and linear attribution methods (including SHAP and Integrated Gradients), showing they can provably fail to improve on random guessing for inferring model behavior on sufficiently rich model classes. Chen et al. (2024) provide the first systematic analysis of the Rashomon effect in gradient boosting specifically, with an information-theoretic characterization of the Rashomon set and a framework for mitigating predictive multiplicity. Marx et al. (2023) demonstrate that SHAP-based feature attributions are sensitive to reference distribution choices. Covert et al. (2021) provide a unified framework connecting removal-based explanations (including SHAP) and discuss stability considerations across methods. The problem is especially acute for correlated features, where SHAP must distribute credit among features that carry overlapping information (Kumar et al., 2020; Chen et al., 2020).
@@
Paillard et al. (2025) argue for computing SHAP values from a single large ensemble rather than aggregating explanations from multiple models, on the grounds that a single ensemble provides a consistent explanation. Our results challenge this recommendation: the single-ensemble approach (our “Ensemble SHAP” baseline) does not resolve the instability caused by correlated features and, when taken to its extreme (the Large Single Model), amplifies it.
Explanation multiplicity.
Hwang et al. (2026) introduce explanation multiplicity— the phenomenon of multiple internally valid but substantively different SHAP explanations for the same decision—and present methodology to disentangle sources due to model training/selection from stochasticity intrinsic to the explanation pipeline. They show that apparent stability depends on the metric: magnitude-based distances can remain near zero while rank-based measures reveal substantial churn in top-feature identity. Our work complements theirs by providing a mechanistic account of the model-training-induced component: first-mover bias explains why retraining gradient-boosted models produces the rank-level instability that Hwang et al. characterize. The FSI diagnostic detects this instability from a single model ensemble without requiring repeated runs. Decker et al. (2024) propose optimal convex combinations of feature attributions to improve robustness or faithfulness; DASH takes a different approach, enforcing diversity at the model level rather than optimizing the aggregation of a fixed set of explanations.
@@ 3.3 Sequential residual dependency
We hypothesize that for a gradient-boosted model $f=\sum_{t=1}^{T}h_{t}$ with $T$ trees, if features $j,k$ belong to a correlated group with pairwise correlation $\rho\to 1$ and tree $h_{1}$ splits on feature $j$, then $\mathbb{E}[|\phi_{j}(f)|]>\mathbb{E}[|\phi_{k}(f)|]$ under TreeSHAP, with the gap increasing in $T$. The expectation is over the randomness in data sampling and split selection. We test the $T$-dependence indirectly in Section 6.1 via the Large Single Model comparison: the LSM uses ${\sim}15{,}000$ sequential trees and produces the poorest reproducibility of any method, consistent with the prediction that more sequential iterations amplify first-mover concentration. Figure 1 provides a direct test: varying $T$ while holding all other factors constant, concentration grows monotonically for a single sequential model and remains flat for $M$ independently trained models averaged at the same total tree budget. Appendix F provides a minimal analytical model of this gain bias.
Line plot showing concentration within a correlated group vs number of trees (n_estimators) for a single sequential model (rising) and an independent ensemble (flat)
Figure 1: Direct isolation of the $T$-scaling prediction. Concentration within a correlated group (max/sum importance, $\rho{=}0.9$, group of 5 features with true importance $0.20$ each) as a function of sequential tree count $T$. Single sequential model (red squares): concentration grows monotonically with $T$, confirming the first-mover prediction. Independent ensemble of $M$ models averaged at matched total tree count (green circles): concentration remains flat regardless of per-model depth. The divergence isolates residual dependency as the operative mechanism. Mean $\pm$ SD across 50 repetitions.
@@ 4.2 Stage 2: Performance Filtering
We retain models whose validation score is within $\varepsilon$ of the best:
$\mathcal{F}=\{i:|s_{i}-s^{*}|\leq\varepsilon\},\quad s^{*}=\max_{i}s_{i},$ (2)
where $s_{i}$ is the negative RMSE (for regression) or AUC (for classification) of model $i$. This ensures that all explanations come from models that have learned meaningful signal. We use $\varepsilon=0.08$ (absolute mode) for synthetic data and $\varepsilon=0.05$ (relative mode, retaining models within 5% of the best score) for real-world datasets.
@@ 4.3 Stage 3: Diversity Selection
This ensures each selected model is maximally different from all previously selected models in its feature utilization pattern, without requiring knowledge of the feature correlation structure. The companion paper (Caraker et al., 2026b) formalizes a balanced ensemble condition: within each collinear group, every feature serves as first-mover equally often, guaranteeing exact within-group equity. Forced low colsample_bytree (Stage 1) and MaxMin selection approximate this balance at smaller $M$ than simple seed averaging, which achieves it only asymptotically via the law of large numbers.
@@ 5.1 Synthetic data
We sweep $\rho\in\{0.0,0.5,0.7,0.9,0.95\}$ with $N_{\text{reps}}=50$ repetitions at each level, regenerating data (same coefficients, new random draws) for each repetition.
@@ 5.2 Real-world datasets
For synthetic data, each of the 50 repetitions regenerates the dataset (same coefficients, new random draws) and retrains all models, capturing both data-sampling and model-selection variance. For real-world datasets, each repetition retrains all models with different random seeds on the same fixed data split, isolating model-selection variance from data-sampling variance. This distinction means that real-world stability estimates reflect only the instability due to arbitrary model choices, not data variability.
@@ 5.3 Methods compared
Independent Stochastic Retrain $K$ models, different seeds, same hyperparameters
@@ 5.4 Evaluation metrics
Mean pairwise Spearman correlation across $N_{\text{reps}}=50$ repeated runs of each method:
@@
BCa bootstrap confidence intervals are computed by resampling the $R$ importance vectors (with replacement), recomputing mean pairwise Spearman $\rho_{S}$ on each bootstrap sample, and applying the bias-corrected and accelerated percentile method.
@@ 6.1 The mechanism: evidence for first-mover bias
Table 3 presents results for the four principal methods across $\rho\in\{0.0,0.5,0.7,0.9,0.95\}$ with 50 repetitions per level. The remaining five baselines (Ensemble SHAP, Single Best M=200M{=}200, LSM Tuned, Random Selection, Naive Top-$N$) are evaluated at $\rho=0.9$ only (Table 4), as their primary purpose is to contextualize the mechanism at high correlation. The key finding is that LSM achieves the poorest stability of any method at every correlation level, despite matching DASH’s total tree count and feature restriction. This is consistent with the first-mover bias hypothesis: sequential residual dependency concentrates importance on arbitrary features, and the effect worsens with correlation severity (LSM stability degrades from $0.955$ at $\rho=0$ to $0.927$ at $\rho=0.95$). Meanwhile, DASH’s stability is effectively flat (0.9730.9730.9770.977), demonstrating immunity to the mechanism it is designed to break.
Table 3: The mechanism experiment: DASH vs. Single Best vs. Large Single Model across correlation levels (50 repetitions per $\rho$). Both DASH and LSM use the same low colsample_bytree; the critical contrast for our hypothesis is that DASH’s models are trained independently while LSM trains trees sequentially. Bold indicates best per metric per $\rho$ level. The four principal methods are shown; all nine are compared at $\rho=0.9$ in Table 4.
@@
0.0 Single Best $972\pm.001 .$985 .163 .611
Large Single Model $955\pm.002 .$976 .170 .772
Stoch. Retrain $\mathbf{.975}\pm.001$ .987 .169 .581
DASH (MaxMin) $973\pm.001 .$985 .153 .606
0.5 Single Best $974\pm.001 .$987 .179 .620
Large Single Model $967\pm.001 .$983 .187 .768
Stoch. Retrain $\mathbf{.979}\pm.001$ .989 .169 .582
DASH (MaxMin) $977\pm.001 .$988 .156 .596
0.7 Single Best $967\pm.002 .$983 .203 .619
Large Single Model $958\pm.002 .$978 .213 .758
Stoch. Retrain $\mathbf{.978}\pm.001$ .989 .178 .583
DASH (MaxMin) $977\pm.001 .$988 .170 .595
0.9 Single Best $958\pm.002 .$978 .232 .613
Large Single Model $938\pm.002 .$968 .258 .733
Stoch. Retrain $\mathbf{.978}\pm.001$ .989 .180 .576
DASH (MaxMin) $977\pm.001 .$988 .175 .591
0.95 Single Best $952\pm.003 .$975 .236 .610
Large Single Model $927\pm.003 .$962 .277 .728
Stoch. Retrain $\mathbf{.980}\pm.002$ .990 .159 .575
DASH (MaxMin) $977\pm.001 .$989 .171 .591
@@
    The effect scales with correlation. LSM stability degrades from $0.955$ to $0.927$ as $\rho$ increases from 0 to $0.95$. DASH stability is flat (0.9730.9730.9770.977). First-mover bias is specifically a collinearity problem.
@@
    Equity degrades in the same pattern. LSM’s within-group CV worsens from $0.170$ to $0.277$, consistent with sequential dependency concentrating importance within correlated groups. DASH distributes it proportionally (0.1530.1530.1750.175).
@@ 6.2 The principle: independence largely resolves it
Table 4 compares all methods at $\rho=0.9$. The critical observation is that DASH and Stochastic Retrain achieve identical stability ($0.977$ each). These methods differ substantially in design— DASH uses forced feature restriction, performance filtering, and diversity-aware selection, while Stochastic Retrain simply trains $K$ models with different random seeds and the same hyperparameters—but they share one property: their explained models are trained independently.
Table 4: All methods at $\rho=0.9$ (50 repetitions), grouped by whether they achieve model independence. Bold indicates best per metric. Matched training budget (M=200M{=}200 models trained). Timing entries marked — share infrastructure with other methods and were not independently measured. 95% BCa bootstrap CIs on stability: Single Best $[0.952,0.963]$, DASH $[0.975,0.978]$ (non-overlapping).
@@
Dependent Single Best $958\pm.002 .$978 .232 1 043.5
Single Best $964\pm.001 .$981 .214 1 248.8
Large Single Model $938\pm.002 .$968 .258 1 006.4
LSM (Tuned) $948\pm.002 .$973 .271 1 088.9
Ensemble SHAP $956\pm.001 .$977 .236 1
Independent Stochastic Retrain $977\pm.002 .$988 .182 30 233.5
Random Selection $976\pm.001 .$988 .187 30 287.2
Naive Top-$N$ $976\pm.001 .$988 .187 30
DASH (MaxMin) $\mathbf{.977}\pm.001$ .988 .176 $\sim$12 140.3
@@
The gap between tiers ($\sim$0.01–0.04) dwarfs the gap within the independent tier (<0.001<0.001). Indeed, Stochastic Retrain achieves marginally higher stability than DASH at every $\rho$ level ($\Delta\approx+0.001$, not significant), confirming that model independence—not DASH’s specific selection mechanism—is the primary driver of stability recovery. DASH’s contributions beyond independence are efficiency ($K_{\text{eff}}\approx 12$ vs. $K=30$ for SR) and equitable attribution spreading (see below).
@@
Within the independent tier, Random Selection ($0.976$) nearly matches DASH ($0.977$), suggesting that MaxMin diversity selection contributes minimally to stability beyond what random sampling from the filtered population already provides. The primary value of diversity selection is therefore not stability but equity: DASH’s within-group CV ($0.176$) is significantly lower than Random Selection’s ($0.187$; Wilcoxon $p<0.001$, Cohen’s d=0.37d=-0.37), confirming that deliberate diversity in feature utilization patterns produces more equitable credit distribution within the filtered-population framework. This equity advantage is consistent across all $\rho$ levels ($p<0.05$ at every level tested). The DASH vs. SR comparison on equity is not significant ($p=0.44$), indicating that the equity benefit is specific to diversity-aware selection, not to the overall pipeline design relative to seed averaging. Practitioners who need only stability can use random selection from a performance-filtered population; those who also require equitable attributions benefit from MaxMin diversity selection.
@@
To directly quantify the role of model-selection randomness, we conduct a fully crossed $7\times 7$ factorial experiment (7 data seeds $\times$ 7 model seeds = 49 cells) and decompose total variance via two-way ANOVA. For Single Best, model selection accounts for $40.6\%$ of total variance—comparable to the $37.6\%$ from data sampling—confirming that model-selection noise is a dominant instability source. DASH shifts the variance budget decisively: data accounts for $73.6\%$ of total variance while model selection drops to $16.2\%$, a $60\%$ reduction in the model-selection sum of squares ($0.636\to 0.089$). The residual (interaction + noise) likewise decreases from $21.8\%$ to $10.2\%$. Table 3 reports the full ANOVA decomposition with $F$-statistics. Both data and model effects are highly significant for both methods ($p<10^{-5}$). The key contrast is between methods: Single Best’s model $F=11.2$ slightly exceeds its data $F=10.3$ (model-dominated), while DASH’s data $F=43.4$ dwarfs its model $F=9.5$ (data-dominated). This shift confirms that independence-based averaging cancels the arbitrary feature choices that dominate single-model explanations.
Table 5: Two-way ANOVA for the $7\times 7$ crossed variance decomposition ($\rho=0.9$). Sum of squares are aggregated across $P=50$ features; $F$-statistics use the residual (interaction + noise) as the error term. [^3] Because each cell contains one $P$-dimensional importance vector and the $S$ are summed across features, these $F$-tests are an aggregate summary rather than a per-feature multivariate test. Per-feature ANOVA tables are available in the code repository. Both data and model effects are significant for both methods; the critical contrast is that DASH shifts from model-dominated (SB) to data-dominated variance.
Method Source $S$ df $F$ $p$
Single Best Data 0.588 6 10.3 $1.2\times 10^{-6}$
Model 0.636 6 11.2 $5.1\times 10^{-7}$
Residual 0.341 36
DASH (MaxMin) Data 0.405 6 43.4 $4.9\times 10^{-15}$
Model 0.089 6 09.5 $2.8\times 10^{-6}$
Residual 0.056 36
As a complementary check, we also conduct a marginal decomposition by alternately fixing the data seed or the model seed. When data is fixed and only model seeds vary, Single Best stability is $0.978$ while DASH achieves 0.9950.995—indicating that DASH nearly eliminates model-selection noise. [^4] The marginal decomposition uses $1-\text{stability}$ as a proxy for instability. Since stability is a mean pairwise Spearman correlation rather than a variance, this is approximate and should be interpreted as directional evidence. The crossed ANOVA above provides exact variance fractions.
@@
Table 6 reports Wilcoxon signed-rank tests with Holm–Bonferroni step-down correction on accuracy and equity metrics across all $\rho$ levels. All comparisons between DASH and LSM are significant at every $\rho$ level with large effect sizes (Cohen’s $d>1.4$). DASH vs. Single Best becomes significant at $\rho\geq 0.7$ for both accuracy ($d=+0.98$) and equity (d=1.03d=-1.03). The DASH vs. Stochastic Retrain comparison is not significant on either accuracy ($d=+0.05$) or equity (d=0.13d=-0.13), consistent with the independence principle. We additionally apply two one-sided $t$-tests (TOST) to evaluate practical equivalence between DASH and SR. The equivalence margin is set adaptively as $\delta=\max(0.01,\;0.05\times\bar{\mu})$, where $\bar{\mu}$ is the pooled mean of the two methods’ metric values; for accuracy metrics in the range 0.930.930.980.98 this evaluates to $\delta\approx 0.049$. We justify this margin by the empirical observation that at $\rho=0.9$, accuracy differences smaller than $0.049$ produce identical top-5 feature rankings in over 95% of repetitions. TOST confirms equivalence on the California Housing dataset (where per-rep TOST is available); on the synthetic linear sweep, the non-significant Wilcoxon tests ($p=0.926$ for accuracy, $p=0.44$ for equity) combined with small effect sizes ($d=+0.05$, d=0.13d=-0.13) provide strong evidence of practical equivalence. Note that stability, computed as a single aggregate across all repetition pairs, cannot be subjected to per-repetition Wilcoxon tests; we report BCa bootstrap confidence intervals (Table 4 caption) as an alternative measure of precision for this metric.
Table 6: Wilcoxon signed-rank tests with Holm–Bonferroni step-down correction (50 paired repetitions). Bold $p$-values are significant at adjusted $\alpha=0.05$. Selected comparisons shown from 26 total tests; full results available in the code repository.